OpenStudy (anonymous):

find an equation of the circle with the center (-2,9) and radius 3

OpenStudy (anonymous):

x^2 + y^2 + 4x - 18y + 76 = 0

OpenStudy (anonymous):

do u want to know how i got this?

OpenStudy (anonymous):

(x+2)^2+(y-9)^2=9

OpenStudy (anonymous):

yes could you please explain

OpenStudy (anonymous):

how did u got this jimmyrep

OpenStudy (anonymous):

Multiplied out what I have.

OpenStudy (anonymous):

Luis has just done that for me though he has 72 anstead of my 76

OpenStudy (anonymous):

If the "moving" part has left any questions, I have used: \[(x-x_{0})^2+(y-y_{0})^2=R^2\]

OpenStudy (anonymous):

is mine right? 85 - 9 = 76

OpenStudy (anonymous):

You can move the equacion of a simple Circle: \[x^2+y^2=R^2\] Where R -> ray of the circle, and you obtain: \[(x+2)^2+(y−9)^2=3^2\] Developing the equation: \[(x^2+4x+4)+(y^2−18y+81)=9\] \[x^2+y^2+4x−18y+76=0\] // Thanks to jimmyrep, i had done it wrong before.

OpenStudy (anonymous):

Ok, I get it now! Thank you both so much!!!

OpenStudy (anonymous):

yw